Find value of the force needed to pull a wheel of mass 1000kg with radius 1.0m over the step of 0.4m


1
Expert's answer
2019-11-05T10:51:07-0500
"mg\sqrt{2rh-h^2}=F(r-h)"

"F=mg\frac{\sqrt{2rh-h^2}}{r-h}"

"F=(1000)(9.8)\frac{\sqrt{2(1)(0.4)-0.4^2}}{1-0.4}=13000\ N"